# On a mean-field Pontryagin minimum principle for stochastic optimal control

### v1

@@ 5. Numerical implementation
Numerical implementations of (30) rely on a set of particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$, which evolve according to
(45a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\}),$
(45b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$
for appropriate Hamiltonian $\mathcal{H}$ and subject to boundary conditions
(46) $X_{0}^{(i)}=a_{i}\sim\pi_{0},\qquad P_{T}^{(i)}=-\nabla_{x}f(X_{T}^{(i)}).$
In order to get closed form ODEs, we need to specify $\mathcal{H}$ which involves approximating $\phi_{t}(x)$ as defined by the constraint (27). We discuss a few choices below. These approximations apply equally well to the discounted cost mean-field PMP formulation from Section 4.
@@ 5.1. A regression formulation
The constraint (27) can be approximated via the method of least squares. For example, one could use
(47) $\mathcal{V}_{t}(\gamma_{t})=\frac{1}{2M}\sum_{i=1}^{M}\|P_{t}^{(i)}+\nabla_{x}\tilde{\phi}(X_{t}^{(i)};\gamma_{t})\|^{2}$
for suitable class of functions $\tilde{\phi}(\cdot;\gamma)$ parametrized by $\gamma\in\mathbb{R}^{d_{\gamma}}$. This approach is not strictly variational unless $\Sigma=0$; but it can be applied to any choice of the gauge variable $\beta_{t}(a)$. See Section 5.3 below for a strictly variational approach.
As a specific example, consider
(48) $\beta_{t}^{(i)}=\frac{1}{2}G(X_{t}^{(i)})R^{-1}G(X_{t}^{(i)})(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)})$
together with a bilinear ansatz
(49) $\tilde{\phi}(x;\gamma)=\frac{1}{2}x^{\rm T}Ax+x^{\rm T}c,$
which makes the term depending on $\Sigma$ in (22) vanish. Thus the Hamiltonian (22) gives rise to
(50a) $\mathcal{H}_{t}(\{X_{t}^{(j)}\},\{P_{t}^{(i)}\})$ $=\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}+(P_{t}^{(i)})^{\rm T}b(X_{t}^{(i)})+c(X_{t}^{(i)})\,+\right.$
(50b) $\qquad\qquad\left.\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}-c_{t}-P_{t}^{(i)})\|^{2}_{R}\right\}$
and the equations of motion (45) reduce to
(51a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
(51b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})\|_{R}^{2}\,-$
(51c) $\qquad\qquad\nabla_{x}(G(X_{t}^{(i)})P_{t}^{(i)})^{\rm T}RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})$
with $A_{t}$ and $c_{t}$ determined by
(52) $A_{t}=C_{t}^{px}(C_{t}^{xx})^{-1},\qquad c_{t}=m^{p}_{t}-A_{t}m^{x}_{t}.$
Here $C_{t}^{px}$ and $C_{t}^{xx}$ denote the empirical covariance matrices between $\{P_{t}^{(i)}\}$ and $\{X_{t}^{(i)}\}$, and $\{X_{t}^{(i)}\}$ with itself, respectively. The empirical mean of $\{X_{t}^{(i)}\}$ and $\{P_{t}^{(i)}\}$, respectively, is denoted by $m_{t}^{x}$ and $m_{t}^{p}$, respectively. Note that $A_{t}$ is not symmetric in general; but can be made symmetric via $1/2(A_{t}^{\rm T}+A_{t})$.
Formulation (51) simplifies further in case $G$ does not depend on $x$; i.e.
(53a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-GRG^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
(53b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+A^{\rm T}G^{\rm T}RG^{\rm T}(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)}).$
@@ 5.2. Discrete Schrödinger bridge approximation
(54b) $\qquad\qquad\sum_{i=1}^{M}\left\{G_{t}^{(i)}(\{X_{t}^{(j)}\})+\beta_{t}^{(i)}\cdot P_{t}^{(i)}+c(X_{t}^{(i)})\right\}$
@@
with $\{v_{t}^{(i)}\}$ chosen such that
@@
Here $\delta_{ij}$ denotes the Kronecker delta function; i.e., $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise. The regularization parameter $\varepsilon>0$ needs to be chosen appropriately [28]. A particularly elegant formulation arises from the choice $\varepsilon=\Delta t$ with $\Delta t>0$ the step-size of an Euler discretization of the Hamiltonian equations of motion [8]. A fast iteterative algorithm for computing the coefficients $v_{t}^{(i)}\geq 0$ has been given in [16, 28].
Writing down the associated Hamiltonian equations of motion (45), we obtain a fully variational implementation of the proposed PMP. However, computing gradients of $m_{t}^{(ij)}$ with respect to particle positions $X_{t}^{(k)}$ is non-trivial and we use the simpler approximation
(59a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}-\sum_{j=1}^{M}m_{t}^{(ij}X_{t}^{(j)},$
(59b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\sum_{j=1}^{M}m_{t}^{(ij)}P_{t}^{(j)}.$
The discrete Schrödinger bridge approximation does not require an explicit approximation of $\phi_{t}(x)$, which simplifies numerical implementations. If $\nabla_{x}\phi_{t}(x)$ is needed it can be obtained via (47).
@@ 5.3. Variational approach
A particularly versatile approach is to approximate $\beta_{t}$ and $\phi_{t}$ via a parametric ansatz $\beta_{t}(x)=\tilde{\beta}(x;\gamma)$ and $\phi_{t}(x)=\tilde{\phi}(x;\alpha_{t})$, respectively, and to consider the empirical Hamiltonian
(60a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha)\,=$
@@
(60c) $\qquad\qquad\left.\tilde{\beta}_{t}(X_{t}^{(j)},\gamma)\cdot\nabla_{x}\left(\tilde{\phi}(X_{t}^{(j)},\alpha)-X_{t}^{(j)}\cdot P_{t}^{(j)}\right)+\frac{1}{2}\Sigma:D_{x}^{2}\tilde{\phi}(X_{t}^{(j)};\alpha)\right\}$
@@
(61a) $\dot{X}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
(61b) $-\dot{P}_{t}^{(i)}$ =($=\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
(61c) $(\gamma_{t},\alpha_{t})$ $=\arg\min_{(\gamma,\alpha)}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha).$
Appropriate choices for $\tilde{\beta}$ and $\tilde{\phi}$ can follow the approaches considered, for exmaple, in [15].
@@ 6. Numerical experiments
We present results from two numerical examples. Both cases involve infinite horizon cost functions of the form (3) and controlled ODEs. We employ the initial value formulation (44) and its numerical approximation using discrete Schrödinger bridges as outlined in Section 5.2. In both cases, we add a diffusion (viscosity) regularization in (22) in line with the concept of viscosity solutions to the associated HJB equations and the vanishing viscosity method [3].
6.1. Inverted pendulum
@@
with $\alpha=25$ and set $R=1$ in (4). We use the discount factor $\gamma=1$ in (3). Note that $(0,0)$ is an unstable equilibrium point under $U_{t}\equiv 0$.
In order to implement and, in fact, regularize the mean-field PMP, we add diffusion (viscosity) by setting $\Sigma=0.1I$ in (22). We initialize the particles at the stable equilibrium point $x_{\rm s}=(\pi,0)$ with added Gaussian noise of variance $0.1I$; i.e. $X_{0}^{(i)}\sim{\rm N}(x_{\rm s},0.1I)$. The initial momenta are set to $P_{0}^{(i)}=-2SX_{0}^{(i)}$ with
(64) $S=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
The discrete Schrödinger bridge based formulation from Section 5.2 has been implemented with ensemble size $M=200$ and $\varepsilon=\Delta t$. The resulting evolution equations (44) have been simulated over 8000 Euler time-steps with time-step $\Delta t=0.025$.
@@
[…]
@@
6.2. Lorenz-63
Figure 2. Controlled Lorenz-63 model: Displayed is the ${\rm x}$-component of a single particle trajectory chosen at random from the ensemble $\{X_{t}^{(i)}\}$ and its associated control $U_{t}^{(i)}$ as a function of time.
@@ 6.2. Lorenz-63
(66) $\dot{X}_{t}=f(X_{t})+U_{t},\qquad f(X)=\left(\begin{array}[]{l}\sigma({\rm y}-{\rm x})\\ -{\rm x}{\rm z}+r{\rm x}-{\rm y}\\ {\rm x}{\rm y}-b{\rm z}\end{array}\right),$
in the state variable $X=({\rm x},{\rm y},{\rm z})^{\rm T}\in\mathbb{R}^{3}$ and with parameters $\sigma=10$, $r=28$, and $b=8/3$ [13]. The control problem is to restrict solutions of (66) to ${\rm x}_{t}\geq 0$ and the authors of [12] introduce the running cost
@@
We have implemented the Schrödinger bridge based mean-field PMP formulation with $\varepsilon=0.1$, step-size $\Delta t=0.001$, and $M=100$ particles. We again added artificial diffusion (viscosity) via $\Sigma=0.5I$ in (22). The results can be found in Figures 2 and 3. The impact of the control can be clearly detected. In particular, by enforcing ${\rm x}_{t}\geq 0$, the chaotic nature of the Lorenz-63 model has been eliminated. The initial spiral towards to quasi-periodic attractor is due to the chosen initialization
(68) $X_{0}=(7.8590;7.1136;27.2293)^{\rm T}+\Xi_{0},\qquad\Xi_{0}\sim{\rm N}(0,0.1I),$
@@
Figure 3. Controlled Lorenz-63 model: Displayed is the three-dimensional trajectory of a single particle chosen at random from the controlled ensemble $\{X_{t}^{(i)}\}$.

### v3

@@ 6. Numerical implementation
Numerical implementations of the McKean–Pontryagin formulations rely on a set of particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$, with corresponding evolution equations which approximate the mean-field ODE formulation (30). At this point two principal choices arise; either one discretizes directly at the level of the mean-field ODE or, alternatively, one discretizes at the level of the Hamiltonian (22) first, giving rise to a Hamiltonian ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$ with associated Hamiltonian equations of motion
(63a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\}),$
(63b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$
subject to boundary conditions
(64) $X_{0}^{(i)}=a_{i}\sim\pi_{0},\qquad P_{T}^{(i)}=\nabla_{x}f(X_{T}^{(i)}).$
Below we discuss three different particle-based approximations which primarily differ in the way the constraint (27) is dealt with. These approximations apply equally well to the discounted cost formulation (47) of the McKean–Pontryagin minimum principle from Section 4.
@@ 6.1. A regression formulation
The constraint (27) can be approximated via the method of least squares. For example, one introduces the cost function
(65) $\mathcal{V}_{t}(\eta)=\frac{1}{2M}\sum_{i=1}^{M}\|P_{t}^{(i)}-\tilde{\Phi}(X_{t}^{(i)};\eta)\|^{2}$
for suitable class of functions $\tilde{\Phi}(\cdot;\eta)$ parametrized by $\eta\in\mathbb{R}^{d_{\eta}}$. The approximation $\tilde{\Phi}(\cdot;\eta_{t})$ with $\eta_{t}=\arg\min\mathcal{V}_{t}(\eta)$ is then used in (30) in place of $\nabla_{x}\phi_{t}(x)$ and gives rise to associated equations of motion in terms of the particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$.
Since we have directly approximated the evolution equations (30), the numerical approximation is not of the form (63) for suitable Hamiltonian $\mathcal{H}$. See Section 6.3 below for a strictly variational approach; i.e., where one first derives a particle approximation $\mathcal{H}$ of the Hamiltonian (22) and subsequently (63).
@@ 6.2. Discrete Schrödinger bridge approximation
(66b) $\qquad\qquad\frac{1}{M}\sum_{i=1}^{M}\left\{S_{t}^{(i)}(\{X_{t}^{(j)}\})+\beta_{t}^{(i)}\cdot P_{t}^{(i)}+c(X_{t}^{(i)})\right\}$
@@
with $\{v_{t}^{(i)}\geq 0\}$ chosen such that
@@
Here $\delta_{ij}$ denotes the Kronecker delta function; i.e., $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise. The regularization parameter $\varepsilon>0$ needs to be chosen appropriately [35]. A particularly elegant formulation arises from the choice $\varepsilon=\Delta t$ with $\Delta t>0$ the step-size of an Euler discretization of the Hamiltonian equations of motion [12]. A fast iterative algorithm for computing the coefficients $v_{t}^{(i)}$ has been given in [22, 35].
Writing down the associated Hamiltonian equations of motion (63), we obtain a fully variational implementation of the proposed McKean–Pontryagin minimum principle. However, computing gradients of $m_{t}^{(ij)}$ with respect to particle positions $X_{t}^{(k)}$ is non-trivial and we use the simpler approximation based on the following direct application of the Schrödinger bridge approximation to (30):
(71a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}-\sum_{j=1}^{M}m_{t}^{(ij)}X_{t}^{(j)},$
(71b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}+\sum_{j=1}^{M}m_{t}^{(ij)}P_{t}^{(j)}.$
The discrete Schrödinger bridge approximation does not require an explicit approximation of $\phi_{t}(x)$, which simplifies numerical implementations. We use this approach for the numerical experiments in Sections 7.1 and 7.2.
@@ 6.3. Variational approach
A particularly versatile approach is to approximate $\beta_{t}$ and $\phi_{t}$ via a parametric ansatz $\beta_{t}(x)=\tilde{\beta}(x;\eta_{t})$ and $\nabla_{x}\phi_{t}(x)=\tilde{\Phi}(x;\alpha_{t})$, respectively, and to consider the Hamiltonian
(72a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t})\,=$
@@
(72c) $\qquad\qquad\left.\tilde{\beta}(X_{t}^{(j)},\eta_{t})\cdot\left(\tilde{\Phi}(X_{t}^{(j)},\alpha_{t})-P_{t}^{(j)}\right)+\frac{1}{2}\nabla_{x}\cdot\left(\Sigma\tilde{\Phi}(X_{t}^{(j)};\alpha_{t})\right)\right\}$
@@
(73a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t}),$
(73b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t}),$
(73c) 0 0 $=\nabla_{\eta}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t})$
(73d) 0 0 $=\nabla_{\alpha}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t}).$
Appropriate choices for $\tilde{\beta}$ and $\tilde{\Phi}$ can follow the approaches considered, for example, in [21]. Below, we discuss the case of linear approximations in some detail.
Example 6.1.
Consider the linear ansatz
(74) $\tilde{\beta}(x;\eta)=Bx+d,\qquad\tilde{\Phi}(x;\alpha)=Ax+c$
for suitable matrices $A,B$ and vector $c,d$; i.e., $\eta=(d,B)$ and $\alpha=(c,A)$. Taking variations with respect to these parameters in (72) leads to
(75) $A_{t}=C_{t}^{px}(C_{t}^{xx})^{-1},\qquad c_{t}=\mu^{p}_{t}-A_{t}\mu^{x}_{t}$
and
(76) $B_{t}=-\frac{1}{2}\Sigma(C_{t}^{xx})^{-1},\qquad d_{t}=-B_{t}\mu_{t}^{x},$
respectively. Here $C_{t}^{px}$ and $C_{t}^{xx}$ denote the empirical covariance matrices between $\{P_{t}^{(i)}\}$ and $\{X_{t}^{(i)}\}$, and $\{X_{t}^{(i)}\}$ with itself, respectively. The empirical mean of $\{X_{t}^{(i)}\}$ and $\{P_{t}^{(i)}\}$, respectively, is denoted by $\mu_{t}^{x}$ and $\mu_{t}^{p}$, respectively. Upon substituting these parameters into (72), the following Hamiltonian equations of motion arise from (63):
(77a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}+\frac{1}{2}\Sigma(C_{t}^{xx})^{-1}(X_{t}^{(i)}-\mu_{t}^{x}),$
(77b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}\,+$
(77c) $\qquad\qquad\frac{1}{2}\Sigma(C_{t}^{xx})^{-1}\left(P_{t}^{(i)}-\mu_{t}^{p}-C^{px}_{t}(C_{t}^{xx})^{-1}(X_{t}^{(i)}-\mu_{t}^{x})\right)\,-$
(77d) $\qquad\qquad\frac{1}{2}C_{t}^{px}(C_{t}^{xx})^{-1}\Sigma(C_{t}^{xx})^{-1}(X_{t}^{(i)}-\mu_{t}^{x}).$
Note that derivatives of the parameters with respect to the particle positions and momenta do not appear because of (73c)–(73d). Such a linear approximation of the McKean–Pontryagin minimum principle is of interest for receding horizon stochastic optimal control problems as they arise from NMPC. See Section 7.3 for a numerical illustration.
We close this section by mentioning that the deep learning approaches exploited in [8, 9, 25, 15] can also be extended to our McKean–Pontryagin formulations. In this case, one would utilize global parametric ansatz functions $\beta_{t}(x)=\tilde{\beta}_{t}(x;\eta)$ and $\nabla_{x}\phi_{t}(x)=\tilde{\Phi}_{t}(x;\alpha)$, $t\in[0,T]$, and directly substitute them into the action principle (21).
@@ 7. Numerical experiments
We present results from three numerical experiments. The first two experiments involve infinite horizon cost functions of the form (3) and controlled ODEs. We employ the initial value formulation (47) with the advection term in (47b) approximated using Nadaraya–Watson kernel regression [3]; i.e.,
(78) $\nabla_{x}\phi_{t}(x)\approx\breve{y}_{t}(x):=\frac{\sum_{i=1}^{M}\exp\left(-\frac{1}{2\delta}\|x-X_{t}^{(i)}\|^{2}\right)P_{t}^{(i)}}{\sum_{i=1}^{M}\exp\left(-\frac{1}{2\delta}\|x-X_{t}^{(i)}\|^{2}\right)}$
for suitable parameter $\delta>0$ and
(79) $D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}\approx\frac{1}{\Delta t}\left(\breve{y}_{t}(X_{t}+\Delta t\dot{X}_{t})-\breve{y}_{t}(X_{t})\right).$
The operator $\mathcal{L}_{\beta}$ is approximated numerically using discrete Schrödinger bridges as outlined in Section 6.2. In both cases, we add a diffusion (viscosity) regularization in (22) in line with the concept of viscosity solutions to the associated HJB equations and the vanishing viscosity method [6].
The final experiment returns to the inverted pendulum from the first experiment but now in the context of NMPC and receding horizon optimal control. Here we rely on an implementation of formulation (77).
7.1. Inverted pendulum: Infinite horizon control
@@
with $\alpha=30$ and set $R=1$ in (4). We use the discount factor $\gamma=1.5$ in (3). Note that $(0,0)^{\rm T}$ is an unstable equilibrium point under $U_{t}\equiv 0$.
[…]
The discrete Schrödinger bridge based formulation from Section 6.2 has been implemented with ensemble size $M=200$ and $\varepsilon=2\Delta t$. The resulting evolution equations (47) have been simulated over 8000 Euler time-steps with time-step $\Delta t=0.05$.
@@
(82) $U_{t}^{(i)}=-G(\theta_{t}^{(i)})^{\rm T}P_{t}^{(i)},\qquad G(\theta)=\left(\begin{array}[]{c}0\\ \cos(\theta)\end{array}\right),$
@@
7.2. Lorenz-63: Infinite horizon control
Figure 2. Controlled Lorenz-63 model: Displayed is the three-dimensional trajectory of the particle $\{X_{t}^{(i)}\}$ with $i=1$ over the time interval $t\in[0,100]$. After an initial transient, the trajectory enters a quasi-period orbit.
@@ 7.2. Lorenz-63: Infinite horizon control
(83) $\dot{X}_{t}=b(X_{t})+GU_{t},\qquad b(x)=\left(\begin{array}[]{l}\sigma({\rm y}-{\rm x})\\ -{\rm x}{\rm z}+r{\rm x}-{\rm y}\\ {\rm x}{\rm y}-b{\rm z}\end{array}\right),$
in the state variable $x=({\rm x},{\rm y},{\rm z})^{\rm T}\in\mathbb{R}^{3}$ and with parameters $\sigma=10$, $r=28$, and $b=8/3$ [19]. The scalar-valued control $U_{t}$ acts only on the ${\rm x}$-component; i.e. $G=(1,0,0)^{\rm T}$. The control problem is to restrict solutions of (83) to ${\rm x}_{t}\geq 0$ and the authors of [18] introduce the running cost
@@
We have implemented the Schrödinger bridge based McKean–Pontryagin formulation with $\varepsilon=0.1$, step-size $\Delta t=0.001$, and $M=100$ particles. We again added artificial diffusion (viscosity) via $\Sigma=0.5I$ in (22). The parameter $\delta$ in (78) is set to $\delta=1$. The initial momenta are $P_{0}^{(i)}=0$.
The results can be found in Figures 3 and 2. The impact of the control can be clearly detected. In particular, by enforcing ${\rm x}_{t}\geq 0$, the chaotic nature of the Lorenz-63 model has been eliminated. The initial spiral towards to quasi-periodic attractor is due to the chosen initialization
(85) $X_{0}^{(i)}=(7.8590,7.1136,27.2293)^{\rm T}+\Xi_{0}^{(i)},\qquad\Xi_{0}^{(i)}\sim{\rm N}(0,0.1I),$
@@
Figure 3. Controlled Lorenz-63 model: Displayed is the dependence of the computed controls $U_{t_{\rm end}}^{(i)}$ as a function of the three components of $X_{t_{\rm end}}^{(i)}$ at final time $t_{\rm end}=100$ for $i=1,\ldots,M$, $M=100$.
7.3. Inverted pendulum: Receding horizon control
We return to the inverted pendulum formulation (80) from Section 7.1 but now in the receding horizon setting of NMPC. A control is computed over receding windows of length $T=0.2$ by implementing (77) with running cost given by (81) with $\alpha=30$ and terminal cost
(86) $f(x)=\frac{\alpha_{T}}{2}\theta^{2}$
with $\alpha_{T}=200$. The computed closed loop control of the form
(87) $u_{t}(x)=-G(\theta)^{\rm T}(A_{t}x+c_{t});$
compare (82) and (75), is then applied to (80) over a prediction interval of length $\Delta\tau=0.02$ with diffusion (Brownian motion) added and $\Sigma=0.01I$. All equations are integrated numerically with step-size $\Delta t=0.002$ and ensemble size $M=100$. The initial ensemble is drawn from the Gaussian ${\rm N}((1,0)^{\rm T},0.01I)$.
We initialize each solution of (77) with the positions from the last prediction step and the momenta from the previous optimization horizon and then solve (77) forward and backward once over a time window of length $T=0.2$. Numerical results can be found in Figure 4. We find that the receding horizon control formulation of the McKean–Pontryagin formulation is able to stabilize the unstable equilibrium $(0,0)^{\rm T}$. The remaining fluctuations in the ensemble are due to the non-vanishing diffusion which guarantee an exploration of state space in the vicinity of $(0,0)^{\rm T}$. Again an optimal choice of $\Sigma$ will depend on the particular problem formulation.
It should be noted that the variational formulation (77) leads to local and linear approximations of the control while the approach used in Section 7.1 delivers globally valid and nonlinear controls for the inverted pendulum. At the same time, the local approximation is easier to compute and suits NMPC.
Figure 4. NMPC for inverted pendulum: Displayed are the time evolution of the ensemble of angles and velocities as a function of time. Starting from an ensemble centered at $(1,0)^{\rm T}$, the controlled ensemble eventually samples the vicinity of the unstable equilibrium $(0,0)^{\rm T}$. The magnitude of the fluctuations depends on the magnitude of the added diffusivity. Those fluctuations decrease as $\Sigma\to 0$.